# Lips Are Lying: Spotting the Temporal Inconsistency between Audio and Visual in Lip-Syncing DeepFakes

### v1

@@ 4.2 Region Awareness
LipSync tends to concentrate on the lower half of the face, and relying solely on coarse-grained global features is insufficient for representation. Therefore, we employ local features to better capture forgery traces.
@@
$\small{\omega_{c_{j}^{i}}=RA([F_{G}|\{F_{R}\}_{j}^{i}],\theta_{RA}),\ c_{j}\in\{c_{h},c_{f},c_{l}\}}$ (3)
where $c_{j}^{i}$ denotes the i-th feature in $c_{j}$ and $\theta_{RA}$ is the parameters of region awareness module. The final feature $F$ is:
$\small{F=\frac{1}{T}\cdot\frac{\sum_{i,j}(\omega_{c_{j}^{i}}\cdot[F_{G}|\{F_{R}\}_{j}^{i}])}{\sum_{i,j}{\omega_{c_{j}^{i}}}}}$ (4)
Region Awareness Loss. We noticed that regardless of what high-level patterns the model learned, it is the lips that count. Other information extracted from the spectrogram and head pose should be served as auxiliary. Hence, we designed $\mathcal{L}_{RA}$, encouraging the region awareness module to focus more on $c_{l}$. Mathematically, the loss is defined as:
$\small{\mathcal{L}_{RA}(\theta_{GR},\theta_{RA})=\sum_{j=1}^{N}\sum_{i=1}^{T}\frac{k}{\exp(\omega_{max}^{i}-\omega_{h}^{i})}}$ (5)
where $\omega_{max}^{i}$ is the max weight in feature stacks, $\omega_{h}^{i}$ is the none-cropped region. $k$ is a hyper-parameter used to adjust the steepness of the loss. With $\mathcal{L}_{RA}$, we expected the model to pay more attention to forged parts.
@@ 4.3 Lip Forgery Detection
Classification. We implement a multi-layer perceptron Taud and Mas (2018) as our classifier optimized with a Binary Cross-Entropy loss:
$\small{\mathcal{L}_{cls}=-\left(y\log(F)+(1-y)\log(1-F)\right)}$ (6)
@@
$\small{\min_{\theta_{RA},\theta_{cls},\theta_{GR}}\omega\cdot\mathcal{L}_{RA}(\theta_{GR},\theta_{RA})+\mathcal{L}_{cls}(\theta_{cls})}$ (7)
@@ 5.1 Setup
Datasets. We trained our model on Wav2Lip-modified LRS2, a subset of our proposed AVLips. We evaluated our method performance on the following datasets: (1) FF++ Rossler et al. (2019), which contains 2,000 samples. (2) DFDC Dolhansky et al. (2020), which has 500 samples. (3) AVLips, our proposed dataset, which includes more than 60,000 samples. Since the baselines we compared against were primarily trained on the FF++ or DFDC datasets, to ensure fairness in the evaluation, we regenerated synthetic data for the first two datasets during the testing phase. This approach aims to maintain consistency and provide a level playing field for a fair comparison of the results.
@@
Metrics. Following existing works Chai et al. (2020); Ojha et al. (2023); Dong et al. (2022), we adopt four popular metrics to get a comprehensive performance evaluation of LipFD. Specifically, we report ACC (accuracy), AP (average precision), FPR (false positive rate), and FNR (false negative rate). We use the AUC (area under the curve) as a metric to evaluate the performance in tackling various perturbation attacks.
@@
Baselines. We take the SOTA methods in general DeepFake detection and lip-based detection as baselines. (1) For image-based DeepFake detections, UniversalDetect Ojha et al. (2023) and SelfBlendedImages Shiohara and Yamasaki (2022) are selected. (2) For video-based DeepFake detections, CViT Wodajo and Atnafu (2021), DoubleStream Shuai et al. (2023) and RealForensics Haliassos et al. (2022) are considered. (3) For the lip-based detection method, we employ the latest LipForensics Haliassos et al. (2021).
5.2 Effectiveness Evaluation
Table 1 shows the performance of LipFD and prior works. We take the most advanced general DeepFake detection methods SelfBlendedImages and UniversalFakeDetect, along with representative DoubleStream and CViT video stream detection models as the baseline for DeepFake detection. We also compared our method with the SOTA lip-based method, namely LipForensics, which guides facial judgment through lip pre-reading. In addition, we have also compared the SOTA multi-modal detection method, namely RealForensics.
Experimental results demonstrate that LipFD outperforms all competitors to a significant extent with a high detection rate and low false alarm rate in detecting the three DeepFake datasets. Also, we find that LipFD attains a commendable precision, as evident from the AP metric. Furthermore, we observe some discernible patterns from Table 1.
First, advanced manipulations are hard to detect by general methods such as UniversalFakeDetect and SelfBlendedImages, indicating that single-frame-based detectors cannot capture dynamic forgeries. In addition, compared to the SOTA RealForensics method, our ACC exceeded it by 3.8%, 88.5%, and 2.1%, respectively. Similar improvements are reflected in the AP as well. This illustrates that concentrating on lip-syncing allows for the extraction of more potential discriminative features than solely observing lip-based movement.
We observe some bad cases from table 1. For example, the AP score is 13% lower than DoubleStream on the FF++ dataset. On the DFDC dataset, our method has an AP lower than LipForensics and RealForensics by 11% and 16%, respectively. As an explanation, these methods primarily aimed at detecting large-scale manipulations of faces in these types of forgeries. In contrast, LipFD focuses on subtle changes in lip inconsistency. Despite a subtle decrease in balance, LipFD still achieves optimal performance in terms of accuracy.
LRS2 (Ours) FF++ DFDC
Method ACC $\uparrow$ AP $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ ACC $\uparrow$ AP $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ ACC $\uparrow$ AP $\uparrow$ FPR $\downarrow$ FNR $\downarrow$
CViT [arXiv 2021] 65.54 56.68 0.07 0.61 62.86 54.17 0.24 0.50 70.99 58.06 0.06 0.50
DoubleStream [IWDW 2023] 75.52 67.72 0.13 0.36 91.02 87.64 0.03 0.14 77.39 69.28 0.21 0.24
UniversalFakeDetect [CVPR 2023] 50.03 50.02 0.99 0.01 50.43 50.16 0.99 0.01 49.86 49.94 0.98 0.01
SelfBlendedImages [CVPR 2022] 49.99 52.13 0.07 0.51 64.59 57.93 0.17 0.53 48.47 49.06 0.15 0.50
RealForensics [CVPR 2022] 91.78 90.14 0.02 0.14 50.43 50.16 0.99 0.01 92.54 91.62 0.01 0.14
LipForensics [CVPR 2021] 86.13 81.56 0.18 0.10 51.02 53.48 0.12 0.46 90.75 87.32 0.08 0.11
@@ 5.2 Effectiveness Evaluation
Table 1: Effectiveness in identifying diverse DeepFakes. Results on LRS2, FF++, and DFDC are reported, including acc, ap, fpr, and fnr. The best result is highlighted in bold, while the second-ranking one is underscored. Throughout the entire experiment, the threshold for the AP metric was set to 0.5.
5.3 Generalizability to Unseen Forgery
A qualified detector is expected to recognize fake videos that were generated using methods not seen during training. In this section, we are going to analyze our model’s generalizability by following the protocol used in Chai et al. (2020); Ojha et al. (2023); Dong et al. (2022).
Method ACC $\uparrow$ AP $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ AUC $\uparrow$
Wav2Lip (train, dynamic) 95.27 93.08 0.04 0.04 95.27
MakeItTalk (static) 96.93 95.49 0.02 0.03 96.89
TalkLip (dynamic) 79.33 80.05 0.34 0.04 80.36
@@ 5.3 Generalizability to Unseen Forgery
5.4 Robustness Evaluation
Robustness analysis aims to evaluate the capability of detectors to withstand common perturbation attacks, as corruptive manipulations on videos are prevalent in the wild, especially in the case of forged videos. Following the setup of RealForensics Haliassos et al. (2022), we train the model on AVLips without data augmentation and then discuss the robustness of the detectors by testing with unseen samples exposed to a set of perturbations. We investigate the performance of the detectors under six types of perturbation at five varying intensities. We use AUC score as evaluation metric, and the experimental results are presented in fig. 6. Evidently, our method outperforms the latest and the best DeepFake detectors RealForensics on most perturbation types.
Our approach effectively against saturation and contrast perturbations which performing linear transformations in the HLS space. For video compression, LipFD exhibits less corruption under varying levels of compression quality. Gaussian blur is applied by maintaining the size of the Gaussian kernel unchanged, while adjusting the standard deviation as the intensity. We find that both blurring and pixelation significantly degrade the performance of the detectors, likely due to the disruption of high-frequency information.
6 Ablation Studies
In this section, we present ablations to understand how our model works. Table 3 shows the overall situation of the experiment. Three significant components, Global feature encoder ($E_{G}$), Global-Region encoder ($E_{GR}$), and Region Awareness module, are ablated from the network separately, and their respective impact on the overall framework was reflected through changes in accuracy metric.
Component ACC $\uparrow$ AP $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ AUC $\uparrow$
@@ 6 Ablation Studies
Full model (basic) 95.27 93.08 0.04 0.04 95.27
Table 3: Overall ablation results. Evaluation scores when components on the left column are stripped away individually.
@@
6.1 Global-Region encoder
Global-Region encoder takes cropped images and a vector encoded by the Global feature encoder as input, merging them into latent codes representing the correlation between regional parts and temporal sequence. The encoder $E_{GR}$ plays a crucial role in the model. As shown in table 3, there is a significant drop in performance when $E_{GR}$ is ablated.
In fig. 8 we visualized the gradients from the last layer of it using Grad-Cam. In the third line with the tag “lip”, the area near the lips has the deepest red color representing the highest gradient. The model focuses almost precisely on the shape of the entire lips. Meanwhile, in the above two lines, the encoder directs its attention to other features, specifically positional information, primarily on the bottom of the heads regardless of real or fake samples, for the reason that LipSync methods predominantly manipulate the bottom parts.
6.2 Region awareness module
The feature stack is assigned different weights by the modules based on their contributions to the discriminator. Subsequently, these features are fused to form the final feature vector. Higher weights indicate a dominant influence during the final judgment. We randomly selected some representative weights, which are shown in fig. 9. All weights are normalized according to the following formula:
$\small{\omega_{i}=\frac{w_{i}}{\omega_{h}+\omega_{f}+\omega_{l}},\ \omega_{i}\in\{\omega_{h},\omega_{f},\omega_{l}\}}$ (8)
For the majority of forged video clips, the crops tagged as “lip” are assigned significantly higher weights than other regions, indicating that these parts contain the most crucial contextual information for discrimination. On the contrary, our module leverages more information in larger-scale images (“face” and “head”) to form the latent code.
Figure 9: The weights assigned by the module dynamically. A higher numerical value indicates that the corresponding region of the image has a more significant impact on the final feature vector.
7 Performance in Real Scenarios
With the advancement of LipSync, certain forgery techniques have been employed for fraudulent purposes. To assess the practicality of our model in real-world scenarios, we conducted experiments across diverse network environments. Our model achieved up to 90.18% accuracy in a network with latency below 100ms which is the common situation of daily life. Results are shown in fig. 7. For more details, please refer to our supplementary materials.
@@ 8 Conclusion
9 Technical Appendix
In the technical appendix, we present more details of our dataset AVLips, robustness evaluation, performance in real-scenario, discussion and future work.
@@ Appendix A AVLips Dataset
In this section, we will provide a detailed description of the features.
@@ A.1 Features
Dynamic expansion. The raw dataset consists of video files in MP4 format and audio files in WAV format. The videos are manipulated using state-of-the-art LipSync generation methods to forge lip movements. The original dataset can be dynamically expanded up to 60 times its initial size using the provided preprocessing code. The expanded samples are represented in the format shown in fig. 10. The randomness algorithm ensures that each expansion generates unique data, ensuring data diversity and providing a convenient data processing approach for temporal detection methods.
@@
Appendix B Robustness Evaluation

### v2

@@ 4.2 Region Awareness
LipSync tends to concentrate on the lower half of the face. Relying solely on coarse-grained global features is insufficient for representation. Hence, we use local features to better capture forgery traces.
@@
$\omega_{c_{j}^{i}}=RA([F_{G}|\{F_{R}\}_{j}^{i}];\theta_{RA}),\ c_{j}\in\{c_{h},c_{f},c_{l}\}$ (3)
where $c_{j}^{i}$ denotes the i-th feature in $c_{j}$ and $\theta_{RA}$ is the parameters of region awareness module $RA(\cdot)$. The final feature $F$ is obtained by concatenating the global feature $F_{G}$ with three series of region features $F_{R}$, which represent the relation between temporal features and region visual features:
$F=\frac{1}{T}\cdot\frac{\sum_{i,j}(\omega_{c_{j}^{i}}\cdot[F_{G}|\{F_{R}\}_{j}^{i}])}{\sum_{i,j}{\omega_{c_{j}^{i}}}}$ (4)
Region Awareness Loss. We noticed that, regardless of the high-level patterns learned by the model, it is the lower part of the face that matters most guan2023stylesync ; ki2023stylelipsync . Other extracted information should be served as auxiliary. Hence, we designed $\mathcal{L}_{RA}$, encouraging the region awareness module to focus more on areas that are more frequently modified. Mathematically, the loss is defined as:
$\mathcal{L}_{RA}=\sum_{j=1}^{N}\sum_{i=1}^{T}\frac{k}{\exp([\omega_{j}^{i}]_{max}-[\omega_{j}^{i}]_{h})}$ (5)
where $\omega_{max}^{i}$ is the max weight in feature stacks, $\omega_{h}^{i}$ is the none-cropped region. $k$ is a hyper-parameter used to adjust the steepness of the loss. With $\mathcal{L}_{RA}$, we hope the model can focus on areas with a higher probability of being modified, such as the face and lips.
@@ 4.3 Lip Forgery Detection
Classification. We implement a multi-layer perceptron mlp as our classifier optimized with a Binary Cross-Entropy loss:
$\mathcal{L}_{cls}=-\left(y\log(F)+(1-y)\log(1-F)\right)$ (6)
@@
$\min_{\theta_{RA},\theta_{cls},\theta_{GR}}\omega\cdot\mathcal{L}_{RA}(\theta_{GR},\theta_{RA})+\mathcal{L}_{cls}(\theta_{cls})$ (7)
@@ 5.1 Setup
Datasets. We trained our model on Wav2Lip-modified LRS3, a subset of our proposed AVLips. We evaluated our method performance on the following datasets: (1) FF++ ff++ , which contains 2,000 samples. (2) DFDC dfdc , which has 500 samples. (3) AVLips, our proposed dataset, which includes more than 20,000 samples. Since the baselines we compared against were primarily trained on the FF++ or DFDC datasets, to ensure fairness in the evaluation, we regenerated synthetic data for the first two datasets during the testing phase. This approach aims to maintain consistency and provide a level playing field for a fair comparison of the results.
@@
Metrics. Following existing works chai2020makes ; UniversalFakeDetect ; caddm , we adopt four popular metrics to get a comprehensive performance evaluation of LipFD. Specifically, we report ACC (accuracy), AP (average precision), FPR (false positive rate), and FNR (false negative rate). We use the AUC (area under the curve) as a metric to evaluate the performance in tackling various perturbation attacks.
@@
Baselines. We take the SOTA methods in general DeepFake detection and lip-based detection as baselines. (1) For image-based DeepFake detections, UniversalDetect UniversalFakeDetect , DoubleStream shuai2023locate and SelfBlendedImages shiohara2022detecting are selected. (2) For video-based DeepFake detections, CViT wodajo2021deepfake and RealForensics haliassos2022leveraging are considered. (3) For the lip-based detection method, we employ the latest LipForensics lipforensic . Detailed information can be found at our Appendix. B.
Table 1: Cross-datasets validation. Results on AVLips, FF++, and DFDC are reported, including acc, ap, fpr, and fnr. The best result is highlighted in bold, while the second-ranking one is underscored. Throughout the entire experiment, the threshold for the AP metric was set to 0.5.
AVLips FF++ DFDC
Method ACC AP FPR FNR ACC AP FPR FNR ACC AP FPR FNR
CViT 65.54 56.68 0.07 0.61 62.86 54.17 0.24 0.50 70.99 58.06 0.06 0.50
DoubleStream 75.52 67.72 0.13 0.36 91.02 87.64 0.03 0.14 77.39 69.28 0.21 0.24
UniversalFakeDetect 50.03 50.02 0.99 0.01 50.43 50.16 0.99 0.01 49.86 49.94 0.98 0.01
SelfBlendedImages 49.99 52.13 0.07 0.51 64.59 57.93 0.17 0.53 48.47 49.06 0.15 0.50
RealForensics 91.78 90.14 0.02 0.14 93.57 91.32 0.03 0.10 92.54 91.62 0.00 0.15
LipForensics 86.13 81.56 0.18 0.10 94.03 93.25 0.04 0.08 90.75 87.32 0.08 0.11
@@
Method ACC FPR FNR AUC
Wav2Lip (Dynamic) 95.27 0.04 0.04 95.27
MakeItTalk (Static) 96.93 0.02 0.03 96.89
TalkLip (Dynamic) 79.33 0.34 0.04 80.36
Table 3: Overall ablation results regarding core modules. We evaluated our model’s performance after removing components listed in the left column.
Component ACC AP FPR FNR AUC
@@
Full model 95.27 93.08 0.04 0.04 95.27
5.2 Effectiveness Evaluation
Table 1 shows the performance of LipFD and prior works. We take the most advanced general DeepFake detection methods SelfBlendedImages and UniversalFakeDetect, along with representative DoubleStream and CViT video stream detection models as the baseline for DeepFake detection. We also compared our method with the SOTA lip-based method, namely LipForensics, which guides facial judgment through lip pre-reading. In addition, we have also compared the SOTA multi-modal detection method, namely RealForensics. Experimental results demonstrate that LipFD outperforms all competitors to a significant extent with a high detection rate and low false alarm rate in detecting the three DeepFake datasets. Also, we find that LipFD attains a commendable precision, as evident from the AP metric. Furthermore, we observe some discernible patterns from Table 1.
First, advanced manipulations are hard to detect by general methods such as UniversalFakeDetect and SelfBlendedImages, indicating that single-frame-based detectors cannot capture dynamic forgeries. In addition, compared to the SOTA RealForensics method, our ACC exceeded it by 3.49%, 1.7%, and 2.73%, respectively. Similar improvements are reflected in the AP as well. This illustrates that concentrating on lip-syncing allows for the extraction of more potential discriminative features than solely observing lip-based movement.
We observe some bad cases from Table 1. For example, the AP score is 16.27% lower than LipForensics on the FF++ dataset. On the DFDC dataset, our method has an AP lower than LipForensics and RealForensics by 13.01% and 12.14%, respectively. As an explanation, these methods primarily aimed at detecting large-scale manipulations of faces in these types of forgeries. In contrast, LipFD focuses on subtle changes in lip inconsistency. Despite a subtle decrease in balance, LipFD still achieves optimal performance in terms of accuracy.
5.3 Generalizability to Unseen Forgery
A qualified detector should recognize fake videos generated by unseen methods. We analyze our model’s generalizability using the protocol from chai2020makes ; UniversalFakeDetect ; caddm .
5.4 Robustness Evaluation
Robustness analysis aims to evaluate the capability of detectors to withstand common perturbation attacks, as corruptive manipulations on videos are prevalent in the wild, especially in the case of forged videos. Following the setup of RealForensics haliassos2022leveraging , we train the model on AVLips without data augmentation and then discuss the robustness of the detectors by testing with unseen samples exposed to a set of perturbations. We investigate the performance of the detectors under six types of perturbation at five varying intensities. We use AUC score as evaluation metric, and the experimental results are presented in Fig. 5. Evidently, our method outperforms the latest and the best DeepFake detectors RealForensics on most perturbation types.
Our approach effectively against saturation and contrast perturbations which performing linear transformations in the HLS space. For compression, LipFD exhibits less corruption under varying levels of quality. Gaussian blur is applied with a fixed kernel size, adjusting the standard deviation for intensity. Both blurring and pixelation significantly degrade detector performance by disrupting high-frequency information.
5.5 Performance in Real Scenarios
With the advancement of LipSync, certain forgery techniques have been employed for fraudulent purposes. To assess the practicality of our model in real-world scenarios, we conducted experiments across diverse network environments. Our model achieved up to 90.18% accuracy in a network with latency below 100ms which is the common situation of daily life zoom2024 ; microsoft2021 ; ookla2024 . Results are shown in Fig. 6. For more details, please refer to our Appendix. D.
6 Ablation Studies
6.1 Core Modules
Table 3 shows the overall situation of the experiment. Three significant components, Global feature encoder ($E_{G}$), Global-Region encoder ($E_{GR}$), and Region Awareness module, are ablated from the network separately, and their respective impact on the overall framework was reflected through changes in accuracy metric.
@@ 6.1 Core Modules
Figure 8: The weights assigned by Region Awareness. A higher value indicates that the corresponding region of the image has a more significant impact on the final feature vector.
Global-Region encoder. Global-Region encoder takes cropped images and a vector encoded by the Global feature encoder as input, merging them into latent codes representing the correlation between regional parts and temporal sequence. The encoder $E_{GR}$ plays a crucial role in the model. As shown in Table 3, there is a significant drop in performance when $E_{GR}$ is ablated. In Fig. 8 we visualized the gradients from the last layer of it using Grad-Cam. In the third line with tag ‘lip’, the area near the lip has the deepest red color representing the highest gradient. The model focuses precisely on the shape of the entire lips. Meanwhile, in the above two lines, the encoder directs its attention to other features, specifically positional information, primarily on the bottom of the heads regardless of real or fake samples, for the reason that LipSync methods predominantly manipulate the bottom parts.
Region awareness. The module assign different weights to the feature stack based on their contributions to discriminator. These features are then fused into the final feature, with higher weights indicating greater influence. Representative weights are shown in Fig. 8, normalized as follows:
$\omega_{i}=\frac{w_{i}}{\omega_{h}+\omega_{f}+\omega_{l}},\ \omega_{i}\in\{\omega_{h},\omega_{f},\omega_{l}\}$ (8)
For the majority of forged video clips, the crops tagged as ‘lip’ are assigned significantly higher weights than other regions, indicating that these parts contain the most crucial contextual information for discrimination. On the contrary, our module leverages more information in larger-scale images (‘face’ and ‘head’) to form the latent code.
6.2 Selection of the Vision Transformers
In this section, we give a comprehensive view of the selection of the ViTs regarding different pretrained datasets and structures. Results are demonstrated in Table 4.
Table 4: Performance under different ViT structures. We selected six popular vision transformers as Global Feature Encoder and tested the final performance of our model.
ViTs ACC AP FPR FNR
CLIP:ViT/L14 95.27 93.08 0.04 0.04
CLIP:ViT/B16 95.00 92.05 0.03 0.07
ImageNet:ViT/L16 93.28 91.13 0.09 0.04
ImageNet:ViT/B16 93.27 91.13 0.09 0.04
ImageNet:Swin-B 94.66 92.53 0.06 0.04
ImageNet:Swin-S 94.59 90.71 0.02 0.09
With the same architecture, the parameter count has a relatively small impact on final performance. Larger pretrained datasets and more challenging pretraining tasks lead to superior model performance and more balanced recognition capabilities (reflected in small differences in False Positive Rate and False Negative Rate). This aligns with our statement in the paper: ‘To effectively carry out its task (capture temporal features), the encoder necessitates extraordinary representational capacity, which can be attained through exposure to a vast number of images’
Under the same pretrained dataset, more advanced model architectures typically lead to better final performances. For example, Swin Transformer achieves better results than vanilla ViTs. This is possibly because the window-based approach employed by Swin Transformer is more suitable for capturing long-term dependencies in video data, assisting in better identification of temporal features.
@@ 7 Conclusion
8 Acknowledgement
This research was supported in part by the National Key Research and Development Program of China under No.2021YFB3100700, the National Natural Science Foundation of China (NSFC) under Grants No. 62202340, 62372334, the CCF-NSFOCUS ‘Kunpeng’ Research Fund under No. CCF-NSFOCUS 2023005, the Open Foundation of Henan Key Laboratory of Cyberspace Situation Awareness under No. HNTS2022004, Wuhan Knowledge Innovation Program under No. 2022010801020127, the Fundamental Research Funds for the Central Universities under No. 2042023kf0121.
@@ Appendix A AVLips Dataset
In this section, we will provide a detailed description of the features. We have introduced the first audio-visual dataset specifically designed for LipSync detection. The goal of this dataset is to solve the issue of many existing DeepFake datasets lacking audio, while also providing foundational support for the field of lip forgery detection.
@@ A.1 Features
Dynamic expansion. The raw dataset consists of video files in MP4 format and audio files in WAV format. The videos are manipulated using state-of-the-art LipSync generation methods to forge lip movements. The original dataset can be dynamically expanded up to 50 times its initial size using the provided preprocessing code. The expanded samples are represented in the format shown in Fig. 9. The randomness algorithm ensures that each expansion generates unique data, ensuring data diversity and providing a convenient data processing approach for temporal detection methods.
@@
Appendix B Experiment Setup
To ensure a fair comparison, we have collected the only known method that employs lipreading from a high dimensional semantic viewpoint for enhancing DeepFake detection (LipForensics lipforensic ), alongside top-performing recent DeepFake detectors. In addition, to fairly compare the performance of methods integrating both video and audio signals, we have collected the latest method that addresses video artifacts and audio-visual inconsistencies (RealForensics haliassos2022leveraging ), which currently stands as the most effective audio-assisted detection method based on our research. The two methods above are both pretrained on FaceForensics++. We evaluated them simultaneously on LipSync and DeepFake detection tasks using well-known datasets FF++, DFDC, and our proposed AVLips.
We ensured fair training settings across all baselines, with a batch size set to 32, the optimizer and the learning rate strictly adhering to the original paper’s settings (Adam, 1e-3$\sim$1e-6). Specifically: 1) Since LipForensics did not provide training scripts, we directly used the pre-trained checkpoints and performed inference under the same experimental settings. For other models, we fine-tuned pretrained weights on AVLips to meet the same baseline; 2) SelfBlendedImages, as an image-based DeepFake detector, was trained with random frames extracted from videos, following the original paper’s settings; 3) UnivefsalFakeDetection, also an image DeepFake detection model, we preprocessed video data into single-frame images for training. Featuring a ResNet core with a Vision Transformer for temporal encoder, our model has approximate 310M parameters pretrained model size, which is comparable to the baselines listed in Table 1.
Appendix C Robustness Evaluation